Question:

1. A firm $ 250,000 under a 5 years term loan agreement at an interest rate of 1o% . The repayment schedule calls for 5 equal payments, the first occurring at the end of the of the first year. What is the amount of each annual payment.

2. Syncor borrowed $ 800,000 payable over 5 years, with an interest rate of 9 percent per annum on the unpaid balance. If the loan is to be repaid in 5 equal, end of year payments, what is the total amount of interest paid on this loan ?

3. Zygo borrowed $ 1 million on a term loan to finance some equipment. The loan required 4 end of year payments of $ 348,821. What is the effective interest rate of this loan ?

4. An intuitively pleasing and prudent investment strategy is to hold:

5. A stock with a beta 0f 1.0 will:

6. Calculate the required rate of return for Mercury, Inc.. assuming that (1) investors except a 4.0% rate of inflation in the future, (2) the real risk free rate is 3.0% (3) the market risk premium is 5.0%, (4) Mercury has a beta of 1.00 and (5) its realized rate of return has averaged 15.0% over the last 5 years.
Answer #1

1.
=250000*10%/(1-1/1.1^5)=65949.3701986863

2.
=800000*9%/(1-1/1.09^5)*5-800000=228369.827826979

3.
=RATE(4,-348821,1000000)=14.796%

4.
high beta stocks in a rising market and low beta stocks in a declining market

5.
behave as market move in tandem with market

6.
=4%+3%+5%=12%